# Panoramic Out-of-Distribution Segmentation for Autonomous Driving

### v2

@@ Abstract
Panoramic imaging enables capturing $360^{\circ}$ images with an ultra-wide Field-of-View (FoV) for dense omnidirectional perception, which is critical to applications, such as autonomous driving and augmented reality, etc. However, current panoramic semantic segmentation methods fail to identify outliers, and pinhole Out-of-distribution Segmentation (OoS) models perform unsatisfactorily in the panoramic domain due to background clutter and pixel distortions. To address these issues, we introduce a new task, Panoramic Out-of-distribution Segmentation (PanOoS), with the aim of achieving comprehensive and safe scene understanding. Furthermore, we propose the first solution, POS, which adapts to the characteristics of panoramic images through text-guided prompt distribution learning. Specifically, POS integrates a disentanglement strategy designed to materialize the cross-domain generalization capability of CLIP. The proposed Prompt-based Restoration Attention (PRA) optimizes semantic decoding by prompt guidance and self-adaptive correction, while Bilevel Prompt Distribution Learning (BPDL) refines the manifold of per-pixel mask embeddings via semantic prototype supervision. Besides, to compensate for the scarcity of PanOoS datasets, we establish two benchmarks: DenseOoS, which features diverse outliers in complex environments, and QuadOoS, captured by a quadruped robot with a panoramic annular lens system. Extensive experiments demonstrate superior performance of POS, with AuPRC improving by $34.25\%$ and FPR95 decreasing by $21.42\%$ on DenseOoS, outperforming state-of-the-art pinhole-OoS methods. Moreover, POS achieves leading closed-set segmentation capabilities and advances the development of panoramic understanding. Code and datasets will be available at PanOoS.
@@ I Introduction
Omnidirectional perception technologies play a crucial role in embodied agents, such as autonomous vehicles and mobile robots, facilitating a comprehensive understanding of the surroundings [1, 2]. Especially, panoramic imaging technology, capable of capturing high-quality $360^{\circ}$ images using optical systems, has gained significant attention cross various omnidirectional vision task [3, 4], such as dense visual prediction [5, 6, 7], holistic scene understanding [8, 9, 10], panoramic scene segmentation [11, 12, 13], and occlusion-aware seamless perception [14]. However, existing panoramic scene segmentation models are trained to recognize a predefined set of semantic classes (e.g., road, building, car, pedestrian[15, 16, 17, 18], fail to identify outliers. Due to the broader Field of View (FoV) of panoramic images and the nature of long-tail semantic distribution in unconstrained surroundings, objects that do not belong to the predefined classes (e.g., animals, furniture) frequently appear in real-world driving environments, which greatly exacerbates the problem. Therefore, there is an urgent need for a technology capable of identifying outliers not present during training to avoid potential dangers.
@@
To enhance comprehensive and safe omnidirectional scene understanding and address this challenge, we introduce a new task, termed the Panoramic Out-of-Distribution Segmentation (PanOoS). Compared with pinhole out-of-distribution segmentation (pinhole-OoS), PanOoS offers a broader FoV beyond the pinhole view, for example, ranging from $95^{\circ}$ to $360^{\circ}$ as depicted in Fig. 1, enabling the detection of outliers and even dangers that are not visible from a pinhole perspective, e.g., the tiger. Yet, panoramic images pose unique challenges in visual understanding tasks due to their broader FoV and complex scene content, which introduce pixel distortions [19, 20] and background clutter, complicating semantic interpretation. These characteristics lead to feature distribution shifts and loss of semantic information, which impair the model’s ability to capture low-level features and fine details of the image. Moreover, these issues interfere with foreground object recognition, introduce semantic ambiguity in overlapping or visually similar regions, and complicate the distinction between in-distribution and out-of-distribution areas.
@@
Figure 3: Impact of background clutter and pixel distortions on state-of-the-art pinhole-OoS methods. Existing pinhole-OoS methods, e.g. Mask2Anomaly and RbA, fail in adapting to the background clutter and pixel distortion, resulting in a higher number of false positives. In contrast, the proposed POS effectively enhances the perception of panoramic pixels.
@@
Although there are some effective techniques for panoramic semantic segmentation [12, 21, 22], they are only designed for closed-set segmentation. The inability to handle outliers means that these techniques may miss critical objects, leading to incorrect or incomplete scene understanding. This issue significantly limits their practical applicability, particularly in dynamic and unpredictable environments where outliers are frequent. High costs associated with data annotation have also resulted in a scarcity of datasets specifically designed for training panoramic segmentation. As illustrated in Fig. 2, most panoramic segmentation methods typically rely on techniques such as unsupervised domain adaptation [13, 21, 14] or knowledge adaptation [23, 24, 25] to transfer models from the pinhole to the panoramic domain. However, these approaches are not suitable for PanOoS, where the model must not only adapt to the characteristics of the panoramic domain, but also learn to distinguish between in-distribution regions and outliers while ensuring closed-set segmentation performance. Meanwhile, state-of-the-art OoS methods [26, 27, 28] for pinhole images typically involve first training a mask-based segmentation model, e.g., in Mask2Former style [29], and then computing anomaly scores using various outlier detection techniques during fine-tuning or inference. Nevertheless, existing research works primarily focus on narrow-FoV pinhole images, making them ill-equipped to address the aforementioned challenges introduced by broader-FoV in panoramic images (see Fig. 3).
@@
To tackle the OoS challenge within the panoramic domain, we propose POS, a novel mask-based framework that integrates prompt distribution learning to enhance pixel-level generalization across complex and distorted panoramic visual domains. (1) Compared to purely visual models, vision-language models are typically exposed to a broader array of visual concepts during pre-training [30]. To enhance the pixel-level perception of outliers within panoramic scenes, POS incorporates CLIP [31] into a mask-transformer architecture, leveraging its strong generalization capabilities to better capture scene semantics. Moreover, during outlier exposure fine-tuning, POS explicitly disentangles per-pixel mask decoding from semantic classification, thereby mitigating semantic entanglement and preserving accurate mask representations for subsequent pixel-level decoding. (2) Traditional image-based query decoding methods [29, 26, 27, 28] often struggle to accurately capture semantic information under pixel distortions and background clutter. To address this limitation, we propose a Prompt-based Restoration Attention (PRA), which integrates the text encoder to extract both class-level and distributional prompts semantic information. By adaptively guiding the learning process, the PRA mechanism optimizes semantic extraction, significantly enhancing the accuracy and robustness of semantic understanding, thereby effectively mitigating semantic ambiguity in challenging panoramic scenes. (3) Finally, an innovative Bilevel Prompt Distribution Learning (BPDL) mechanism is proposed. By leveraging the language-image semantic distribution consistency of the CLIP model, BPDL refines the spatial distribution of per-pixel mask feature representations and strengthens the distinguishability of out-of-distribution features and the adaptability of the pinhole-trained model to severe distortions encountered in panoramic images. With these crucial designs, POS achieves superior alignment with the characteristics of panoramic images, enabling efficient out-of-distribution segmentation while demonstrating impressive closed-set segmentation performance.
@@
To facilitate the evaluation of the PanOoS task, we create two fresh datasets: (1) We have spent a large effort on artificial screening and built the first panoramic out-of-distribution segmentation dataset: Dense Panoramic Out-of-distribution Segmentation (DenseOoS), for model evaluation. Due to the broader FoV, the backgrounds of each panoramic image are very complex, featuring object instances that vary in shape and appearance, along with significant distortions. Meanwhile, outliers are randomly distributed across reasonable areas. (2) To further verify the performance of POS in practical application and explore the PanOoS in unconstrained, unstructured, and intense-motion real-world scenarios, we employ a quadruped mobile robot to collect the real motion scene panoramic images with outliers, creating the Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. Compared to DenseOoS, QuadOoS contains a greater number of abnormal objects in each image, with smaller sizes, while maintaining the same panoramic pixel characteristics. Ultimately, based on DenceOoS and QuadOoS, we established the evaluation benchmarks for the PanOoS task. The proposed POS achieves state-of-the-art performance on both two benchmarks, significantly outperforming other pinhole-OoS methods [32, 33, 28, 27, 26], with AuPRC, and FPR95 of $(85.56\%,48.13\%)$, and $(0.45\%,85.22\%)$, respectively. Furthermore, compared to panoramic semantic segmentation methods [34, 12, 2], it demonstrates superior closed-set segmentation performance, with an mIoU of $63.30\%$. These results collectively validate the effectiveness of the proposed POS in addressing PanOoS and enhancing panoramic scene understanding.
@@
  • We introduce the Panoramic Out-of-Distribution Segmentation (PanOoS) task, which, for the first time, explores out-of-distribution segmentation in the panoramic domain, aiming to facilitate comprehensive and holistic scene understanding.
  • To achieve the PanOoS task, we propose POS, a mask-transformer-based model with Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), which effectively alleviates the effects of background clutter and pixel distortions, demonstrating remarkable performance.
@@ II Related Work
Panoramic Semantic Segmentation. By capturing broader-FoV scenes, panoramic images serve as a starting point for a more holistic understanding of $360^{\circ}$ surroundings [11, 35]. The first line of research on panoramic semantic segmentation [36, 37, 22, 38, 39] is based on supervised learning. However, to address the scarcity of annotations, researchers have revisited panoramic segmentation from the perspective of unsupervised domain adaptation [40, 41, 42, 43, 44] by leveraging rich training sources from narrow-FoV pinhole data. The approaches can be divided into three types: pseudo-labeling [45, 8], adversarial training [18, 13, 46], and prototype adaptation [2, 47, 14, 48, 49] methods. Pseudo-labeling generates self-supervised labels for the panoramic domain. Adversarial training enforces the target model to capture domain-invariant characteristics across domains using a discriminator. Prototype adaptation approaches align high-level feature centers between the pinhole and panoramic domains. Additionally, related work demonstrates strong performance from the perspective of knowledge adaptation [23, 25, 24]. However, existing methods focus on closed-set segmentation and do not address the problem of identifying outliers within the scene. Recently, open panoramic segmentation [10] has been proposed, but the focus has been on improving the efficiency of panoramic open-vocabulary segmentation. Different from previous works, we aim to explore out-of-distribution segmentation in panoramic images and exploit the distribution learning of semantic text-based prompts to advance panoramic scene understanding.
@@
Out-of-distribution Segmentation. Out-of-distribution segmentation aims to identify and separate outliers in some specific scenes, such as complex driving environments [15, 50, 51, 52], while maintaining closed-set segmentation capability. Early uncertainty-based methods assume that outliers lead to low-confidence predictions, focusing on estimating the prediction uncertainty through various ways to measure pixel-wise anomaly scores [53, 54, 55, 56, 57]. However, models trained on closed sets may make high-confidence erroneous predictions for unseen categories. Follow-up works have shifted their focus to reconstruction. Reconstruction-based methods detect outliers by comparing the differences between the original image and the reconstruction generated from semantic segmentation predictions [51, 58, 59, 60, 61, 62]. However, nearly all reconstruction-based methods are heavily dependent on the quality of the generated output, which can lead to performance degradation due to image artifacts [63]. These methods also suffer from context dependence and high computational costs, resulting in unsatisfactory performance in complex environments.
@@
Recent approaches [64, 32, 33, 65, 66] introduce the Outlier Exposure (OE) strategy [67] that utilizes auxiliary out-of-distribution data to regularize the model’s feature space. These auxiliary outliers typically are images from other datasets (such as ImageNet [68], COCO [69], ADE20K [70], etc.) or involve cutting out abnormal objects from other images and pasting them into inlier scenes. Notably, PEBAL [32] learns energy-based adaptive penalties through abstention learning, whereas DenseHybrid [33] combines likelihood and posterior evaluation to achieve better results. Compared to earlier works, the OE strategy yields better performance, but pixel-wise architecture methods score each pixel individually without considering local semantics, leading to increased noise in anomaly predictions. Currently, mask-transformer-based methods [71, 26, 28, 27, 72, 73] overcome this limitation by segmenting outliers into semantically clustered masks and encouraging the objectness of predictions, which achieve leading segmentation results. EAM [26] ensembles region-based outlier scores, while Mask2Anomaly [28] combines contrastive learning and mask refinement to significantly reduce false positives. RbA [27] utilizes mask queries that act as a one-vs-all classifier [26, 27, 74], helping to eliminate uncertainty in ambiguous in-point regions, e.g., semantic boundaries, by reducing the scores in negative pixels. However, existing methods primarily focus on narrow FoV pinhole images, and when transferred to the panoramic domain, they struggle to adapt to challenges such as background clutter and pixel distortions caused by the broad FoV. To the best of our knowledge, this is the first work to tackle this pressing problem in panoramic out-of-distribution segmentation.

### v3

@@ Abstract
Panoramic imaging enables capturing $360^{\circ}$ images with an ultra-wide Field-of-View (FoV) for dense omnidirectional perception, which is critical to applications, such as autonomous driving and augmented reality, etc. However, current panoramic semantic segmentation methods fail to identify outliers, and pinhole Out-of-distribution Segmentation (OoS) models perform unsatisfactorily in the panoramic domain due to pixel distortions and background clutter. To address these issues, we introduce a new task, Panoramic Out-of-distribution Segmentation (PanOoS), with the aim of achieving comprehensive and safe scene understanding. Furthermore, we propose the first solution, POS, which adapts to the characteristics of panoramic images through text-guided prompt distribution learning. Specifically, POS integrates a disentanglement strategy designed to materialize the cross-domain generalization capability of CLIP. The proposed Prompt-based Restoration Attention (PRA) optimizes semantic decoding by prompt guidance and self-adaptive correction, while Bilevel Prompt Distribution Learning (BPDL) refines the manifold of per-pixel mask embeddings via semantic prototype supervision. Besides, to compensate for the scarcity of PanOoS datasets, we establish two benchmarks: DenseOoS, which features diverse outliers in complex environments, and QuadOoS, captured by a quadruped robot with a panoramic annular lens system. Extensive experiments demonstrate superior performance of POS, with AuPRC improving by $34.25\%$ and FPR95 decreasing by $21.42\%$ on DenseOoS, outperforming state-of-the-art pinhole-OoS methods. Moreover, POS achieves leading closed-set segmentation capabilities and advances the development of panoramic understanding. Code and datasets will be available at PanOoS.
@@ I Introduction
Omnidirectional perception technologies are vital for embodied agents, such as autonomous vehicles and mobile robots, enabling comprehensive understanding of the surroundings [zhang2024behind, ai2025survey]. In particular, panoramic imaging technology, capable of capturing high-quality $360^{\circ}$ images via optical systems, has gained significant attention across various omnidirectional vision tasks [zheng2025one_flight], such as dense visual prediction [shen2022panoformer, ling2023panoswin, yu2023panelnet], holistic scene understanding [zhang2021deeppanocontext, zheng2025open], and panoramic scene segmentation [yang2020omnisupervised, zhang2022bending, zheng2023both, cao2024occlusion]. However, existing panoramic semantic segmentation models are trained to recognize a predefined set of semantic classes (e.g., road, building, car, pedestrian[cordts2016cityscapes, ma2021densepass, huang2024high], which fail to identify outliers. Due to the broader Field-of-View (FoV) of panoramic images and the long-tail nature of semantic distribution in unconstrained surroundings, objects that do not belong to the predefined classes (e.g., animals, furniture) frequently appear in real-world driving environments, which greatly exacerbates the problem. Therefore, there is an urgent need for a technology capable of identifying outliers not present during training to avoid potential dangers.
@@
To enhance comprehensive and safe omnidirectional scene understanding, we introduce a new task, termed Panoramic Out-of-Distribution Segmentation (PanOoS). Unlike pinhole out-of-distribution segmentation (pinhole-OoS), PanOoS offers a broader FoV beyond the pinhole view, for example, ranging from $95^{\circ}$ to $360^{\circ}$ as depicted in Fig. 1, enabling detection of outliers and even dangers not visible from a pinhole perspective, e.g., a tiger. Yet, panoramic images pose unique challenges in visual understanding tasks due to their broader FoV and complex scene content, which introduce pixel distortions [gao2022review, yang2021capturing] and background clutter, thereby complicating semantic interpretation. These characteristics lead to feature distribution shifts and loss of semantic information, impairing the model’s ability to capture low-level features and fine details of the image. Moreover, these issues interfere with foreground object recognition, introduce semantic ambiguity in overlapping or visually similar regions, and complicate distinction between in-distribution and out-of-distribution areas.
@@
Figure 3: Impact of pixel distortions and background clutter on state-of-the-art pinhole-OoS methods. Existing pinhole-OoS methods, e.g. Mask2Anomaly and RbA, fail to adapt to pixel distortions and background clutter, resulting in a higher number of false positives. In contrast, the proposed POS effectively enhances semantic perception of panoramic pixels.
@@
Although several effective techniques exist for panoramic semantic segmentation [zhang2022bending, zheng2023look, li2023sgat4pass], they are only designed for closed-set segmentation. The inability to handle outliers means that these techniques may miss critical objects, leading to incorrect or incomplete scene understanding. This issue significantly limits their practical applicability, particularly in dynamic and unpredictable environments where outliers frequently occur. High costs associated with data annotation have also resulted in a scarcity of datasets specifically designed for training panoramic segmentation models. As illustrated in Fig. 2, most panoramic segmentation methods rely on techniques such as unsupervised domain adaptation [zheng2023both, zheng2023look, cao2024occlusion] or knowledge adaptation [zhang2024goodsam, jaus2023panoramic_insights, zhang2024goodsam++] to transfer models from the pinhole to the panoramic imagery. However, these approaches are not suitable for PanOoS, where the model must not only adapt to the characteristics of the panoramic imagery, but also learn to distinguish between in-distribution regions and outliers while ensuring closed-set segmentation performance. Meanwhile, state-of-the-art OoS methods [grcic2023advantages, nayal2023rba, rai2024mask2anomaly] for pinhole images typically involve first training a mask-based segmentation model, e.g., in Mask2Former style [cheng2022masked], and then computing anomaly scores using various outlier detection techniques during fine-tuning or inference. Nevertheless, existing research works primarily focus on narrow-FoV pinhole images, and thus lack the architectural or learning mechanisms needed to handle the geometric distortions and wide contextual dependencies inherent in panoramic imagery, making them ill-equipped for PanOoS (see Fig. 3).
@@
To tackle the OoS challenge within the panoramic imagery, we propose POS, a mask-based framework that integrates prompt distribution learning to enable accurate pixel-wise detection of out-of-distribution regions in complex and distorted panoramic scenes. (1) Compared to purely visual models, vision-language models are typically exposed to a broader array of visual concepts during pre-training [yu2023convolutions]. To enhance the pixel-level perception of outliers within panoramic scenes, POS incorporates CLIP [radford2021learning] into a mask-transformer architecture, leveraging its strong generalization capabilities to better capture scene semantics. Moreover, during outlier exposure fine-tuning, POS explicitly disentangles per-pixel mask decoding from semantic classification, thereby mitigating semantic entanglement and preserving accurate mask representations for subsequent pixel-level decoding. (2) Traditional image-based query decoding methods [cheng2022masked, grcic2023advantages, nayal2023rba, rai2024mask2anomaly] often struggle to accurately capture semantic information under pixel distortions and background clutter. To address this, we propose Prompt-based Restoration Attention (PRA), which leverages the text encoder to extract semantic information from prompts at both class-level and distributional levels. By adaptively guiding the learning process, the PRA mechanism optimizes semantic extraction, significantly enhancing the accuracy and robustness of semantic understanding, thereby effectively mitigating semantic ambiguity in challenging panoramic scenes. (3) Finally, an innovative Bilevel Prompt Distribution Learning (BPDL) mechanism is proposed. By leveraging the language-image semantic distribution consistency of the CLIP model, BPDL refines the distribution of per-pixel mask embeddings, strengthens the distinguishability of out-of-distribution features and the adaptability of the pinhole-trained model to severe distortions encountered in panoramic images. With these crucial designs, POS achieves superior alignment with the characteristics of panoramic images, enabling efficient out-of-distribution segmentation while demonstrating impressive closed-set segmentation performance.
@@
To facilitate the evaluation of the PanOoS task, we create two fresh datasets: (1) We have spent significant effort on artificial screening and built the first panoramic out-of-distribution segmentation dataset: Dense Panoramic Out-of-distribution Segmentation (DenseOoS). Due to the broader FoV, the backgrounds of each panoramic image are highly complex, featuring object instances that vary in shape and appearance, along with significant distortions. Meanwhile, outliers are placed in semantically plausible regions to ensure realism. (2) To further verify the performance of POS in practical application and explore the PanOoS in unconstrained, unstructured, and intense-motion real-world scenarios, we employ a quadruped mobile robot to collect panoramic images containing outliers, creating the Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. Compared to DenseOoS, QuadOoS contains a greater number of outliers per image, with smaller sizes, while maintaining similar panoramic imaging characteristics. Ultimately, based on DenseOoS and QuadOoS, we establish the evaluation benchmarks for the PanOoS task. The proposed POS achieves state-of-the-art performance on both benchmarks, significantly outperforming other pinhole-OoS methods [tian2022pixel, grcic2022densehybrid, rai2024mask2anomaly, nayal2023rba, grcic2023advantages], with AuPRC, and FPR95 of $(85.56\%,48.13\%)$, and $(0.45\%,85.22\%)$, respectively. Furthermore, compared to panoramic semantic segmentation methods [hoyer2022daformer, zhang2022bending, zhang2024behind], it demonstrates superior closed-set segmentation performance, with an mIoU of $63.30\%$ on DenseOoS. These results collectively validate the effectiveness of the proposed POS in addressing PanOoS and enhancing panoramic scene understanding.
@@
  • We introduce the Panoramic Out-of-Distribution Segmentation (PanOoS) task, which, for the first time, explores out-of-distribution segmentation in panoramic imagery, aiming to facilitate holistic scene understanding.
  • To achieve the PanOoS task, we propose POS, a mask-transformer-based model with Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), which effectively alleviates the effects of pixel distortions and background clutter, demonstrating remarkable performance.
@@ II-A Panoramic Semantic Segmentation
By capturing broader FoV scenes, panoramic images serve as a starting point for a more holistic understanding of $360^{\circ}$ surroundings [yang2019pass, jiang2024minimalist]. The first line of research on panoramic semantic segmentation [xu2019semantic, orhan2022semantic_outdoor, li2023sgat4pass, kim2022pasts, guttikonda2024single] is based on supervised learning. However, to address the scarcity of annotations, researchers have revisited panoramic segmentation from the perspective of unsupervised domain adaptation [jang2022dada, zhu2023patch, zhang2021transfer, jiang2025gaussian, yuan2023laformer] by leveraging rich training sources from narrow-FoV pinhole data. The approaches can be divided into three types: pseudo-labeling [liu2021pano, zhang2021deeppanocontext], adversarial training [ma2021densepass, zheng2023both, jiang2025multi], and prototype adaptation [zhang2024behind, zheng2024semantics, cao2024occlusion, zheng2024360sfuda++, zhong2025omnisam]. Pseudo-labeling generates self-supervised labels for the panoramic domain. Adversarial training encourages the target model to capture domain-invariant characteristics across domains using a discriminator. Prototype adaptation approaches align high-level feature centers between the pinhole and panoramic domains. Additionally, related work demonstrates strong performance from the perspective of knowledge adaptation [zhang2024goodsam, zhang2024goodsam++, jaus2023panoramic_insights]. However, existing methods focus on closed-set segmentation and do not address the problem of identifying outliers within the scene. Recently, open panoramic segmentation [zheng2025open] has been proposed, but the focus has been on improving the efficiency of open-vocabulary segmentation. In contrast to prior work, we aim to explore out-of-distribution segmentation in panoramic images and leverage the distribution learning of text-based prompts to advance panoramic scene understanding.
II-B Out-of-distribution Segmentation
@@
Out-of-distribution segmentation aims to identify and separate outliers in specific scenes, such as complex driving environments [chan2021segmentmeifyoucan, blum2021fishyscapes, cen2023consensus, wan2025out], while maintaining closed-set segmentation capability. Early uncertainty-based methods assume that outliers lead to low-confidence predictions and focus on estimating the prediction uncertainty through various approaches to measure pixel-wise anomaly scores [mukhoti2018evaluating, corbiere2019addressing, hendrycks2019scaling, jung2021standardized, cen2021deep]. However, models trained on closed sets may make high-confidence erroneous predictions for unseen categories. Subsequent works have shifted their focus to reconstruction. Reconstruction-based methods detect outliers by measuring the discrepancy between the original image and a reconstruction generated from semantic segmentation predictions [lis2019detecting, xia2020synthesize, haldimann2019not, vojir2021road, grcic2021dense, lis2023detecting]. However, nearly all reconstruction-based methods are highly sensitive to the quality of the reconstructed output, which can lead to performance degradation due to image artifacts [fontanel2021detecting]. These methods also suffer from context dependence and high computational overhead, resulting in unsatisfactory performance in complex environments.
@@
Recent approaches [chan2021entropy, tian2022pixel, grcic2022densehybrid, liu2023residual, choi2023balanced] leverage the Outlier Exposure (OE) strategy [hendrycks2018deep] that utilizes auxiliary out-of-distribution data to regularize the model’s feature space. These auxiliary outliers are typically images from other datasets (such as ImageNet [deng2009imagenet], COCO [lin2014microsoft], ADE20K [zhou2017scene], etc.) or synthetic composites created by cutting out anomalous objects from external images and pasting them into inlier scenes. Notably, PEBAL [tian2022pixel] learns energy-based adaptive penalties through abstention learning, whereas DenseHybrid [grcic2022densehybrid] combines likelihood and posterior evaluation to achieve better results. Although the OE strategy generally improves performance over earlier approaches, pixel-wise architectures that adopt it still score each pixel independently without considering local semantic context, leading to noisy anomaly predictions. Currently, mask-transformer-based methods [ackermann2023maskomaly, grcic2023advantages, rai2024mask2anomaly, nayal2023rba, delic2024outlier, zhang2024csl] overcome this limitation by segmenting outliers into semantically clustered masks and encouraging the objectness of predictions, achieving leading segmentation results. EAM [grcic2023advantages] ensembles region-based outlier scores, while Mask2Anomaly [rai2024mask2anomaly] combines contrastive learning and mask refinement to significantly reduce false positives. RbA [nayal2023rba] utilizes mask queries that act as a one-vs-all classifier [grcic2023advantages, nayal2023rba, shoeb2025out], helping to eliminate uncertainty in ambiguous in-point regions, e.g., semantic boundaries, by reducing the scores in negative pixels. However, existing methods primarily focus on narrow-FoV pinhole images, and when transferred to the panoramic domain, they struggle to adapt to challenges such as pixel distortions and background clutter caused by the broad FoV. To the best of our knowledge, this is the first work to tackle this pressing problem of out-of-distribution segmentation in panoramic imagery.